# Rethinking the Capacity of Graph Neural Networks for Branching Strategy

### v2

@@ 1. Introduction
Branching is the process of dividing a linear relaxation into two subproblems. When branching, the solver must select a variable with a fractional value in the relaxation’s solution and create two new subproblems. In one subproblem, the variable is forced to be less than or equal to the nearest integer below the fractional value. In the other, it is forced to be greater than or equal to the nearest integer above. The branching variable choice is critical because it can impact the solver’s efficiency by orders of magnitude.
@@
Learning to branch (L2B) introduces a new approach by incorporating machine learning (ML) to develop branching strategies, offering new solutions to address this trade-off. This line of research begins with imitation learning [khalil2016learning, alvarez2017machine, balcan2018learning, gasse2019exact, gupta2020hybrid, zarpellon2021parameterizing, gupta2022lookback, lin2022learning, yang2022learning], where models, including SVM, decision tree, and neural networks, are trained to mimic SB outcomes based on the features of the underlying MILP. They aim to create a computationally efficient strategy that achieves the effectiveness of SB on specific datasets. Significant progress has been made by training such models on specific MILP datasets and applying the learned strategies to similar problems. Furthermore, in reinforcement learning approaches, mimicking SB continues to take crucial roles in initialization or regularization [qu2022improved, zhang2022deep].
@@
  • In the context of L2B using GNNs, we first focus on the most widely used type: message-passing GNNs (MP-GNNs). Our study reveals that MP-GNNs can reliably predict SB results, but only for a specific class of MILPs that we introduce as message-passing-tractable (MP-tractable). We prove that for any distribution of MP-tractable MILPs, there exists an MP-GNN capable of accurately predicting their SB results. This finding establishes a theoretical basis for the widespread use of MP-GNNs to approximate SB results in current research, validating their applications within MP-tractable MILPs.
@@ 2. Preliminaries and problem setup
$\min~{}~{}\begin{bmatrix}~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1}}&~{}~{}~{}~{}~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{2}}&~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{3}}~{}~{}~{}\end{bmatrix}\begin{bmatrix}1\\ 2\\ 3\end{bmatrix}$s.t.
$2~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1}}$ $+$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{2}}$ $\leq 5$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{2}}$ $+$ $3~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1}}~{}~{}~{},~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{2}}~{}~{}~{}~{}~{},~{}~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{3}}~{}\leq 1$
$x_{1}\in\mathbb{Z}$
$w_{1}$$(1,0,1,1)$$w_{2}$$(2,0,1,0)$$w_{3}$$(3,0,1,0)$$v_{1}$$(5,\leq)$$v_{2}$$(0,\geq)$2113
@@
Graph Representation of MILP
Here we present an approach to represent MILP as a bipartite graph, termed the MILP-graph. This conceptualization was initially proposed by [gasse2019exact] and has quickly become a prevalent model in ML for MILP-related tasks. The MILP-graph is defined as a tuple $G=(V,W,A,F_{V},F_{W})$, where the components are specified as follows:
  • $V=\{1,2,\dots,m\}$ and $W=\{1,2,\dots,n\}$ are sets of nodes representing the constraints and variables, respectively,
  • An edge $(i,j)$ connects node $i\in V$ to node $j\in W$ if the corresponding entry $A_{ij}$ in the coefficient matrix of (2.1) is non-zero, with $A_{ij}$ serving as the edge weight,
  • $F_{V}$ are features/attributes of constraints, with features $v_{i}=(b_{i},\circ_{i})$ attached to node $i\in V$,
  • $F_{W}$ are features/attributes of variables, with features $w_{j}=(c_{j},\ell_{j},u_{j},\delta_{I}(j))$ attached to node $j\in W$, where $\delta_{I}(j)\in\{0,1\}$ indicates whether the variable $x_{j}$ is integer-constrained.
@@
Message-passing graph neural networks (MP-GNNs) are a class of GNNs that operate on graph-structured data, by passing messages between nodes in a graph to aggregate information from their local neighborhoods. In our context, the input is an aforementioned MILP-graph $G=(V,W,A,F_{V},F_{W})$, and each node in $W$ is associated with a real-number output. We use the standard MP-GNNs for MILPs in the literature [gasse2019exact, chen2022representing-milp]. Specifically, the initial layer assigns features $s_{i}^{0},t_{j}^{0}$ for each node as
@@
Any $2$-FGNN [^3] It’s important to note that 2-FGNNs tailored for MILP-graphs differ from those designed for general graphs. Specifically, in the MILP-graph context, there’s no requirement to consider pairs of constraints (such as $(i_{1},i_{2})$ where $i_{1},i_{2}\in V$.). This distinction is underscored by Theorem 4.7 discussed later., $F\in\mathcal{F}_{\textup{2-FGNN}}$, maps a MILP-graph $G$ to a $n$-dim vector: $y=F(G)$. While MP-GNNs and 2-FGNNs share the same input-output structure, their internal structures differ, leading to distinct expressive powers.
@@ 3. Imitating strong branching by GNNs
Observation I: SB is notably effective in reducing the size of the BnB search space. This size is measured by the size of the BnB tree. Here, a “tree” refers to a hierarchical structure of “nodes”, each representing a decision point or a subdivision of the problem. The tree’s size corresponds to the number of these nodes. For instance, consider the instance “neos-3761878-oglio” from MIPLIB [gleixner2021miplib]. When solved using SCIP [BolusaniEtal2024OO, BolusaniEtal2024ZR] under standard configurations, the BnB tree size is $851$, and it takes $61.04$ seconds to attain optimality. However, disabling SB, along with all branching rules dependent on SB, results in an increased BnB tree size to $35548$ and an increased runtime to $531.0$ seconds.
Observation II: SB itself is computationally expensive. In the above experiment under standard settings, SB consumes an average of $70.40\%$ of the total runtime, $42.97$ out of $61.04$ seconds in total.
@@
We give several remarks for Definition 3.2. We only define the SB score for MILP problems with feasible and bounded LP relaxations; otherwise the optimal solution $x^{*}_{\textup{LP}}(G)$ does not exist. If the LP relaxation of $G$ admits multiple optimal solutions, then the strong branching score $\textup{SB}(G)$ depends on the choice of the particular optimal solution. To guarantee that the SB score is uniquely defined, in Definition 3.2, we use the optimal solution with the smallest $\ell_{2}$-norm, which is unique.
@@ 4. Main results
This section presents our main results for the expressive power of MP-GNNs and 2-FGNNs to represent SB scores of MILP problems.
@@ 4.1. MP-GNNs can represent SB for MP-tractable MILPs
In this section, we define a class of MILPs, named message-passing-tractable (MP-tractable) MILPs, and then establish an universal approximation result for MP-GNNs to represent SB within this class.
To define MP-tractability, we first present the Weisfeiler-Lehman (WL) test [weisfeiler1968reduction], as it is well-known in the literature that the WL test is a pivotal criterion for assessing the expressive power of MP-GNNs [xu2019powerful]. The WL test in the context of MILP-graphs is stated in Algorithm 1. It follows exactly the same updating rule as the MP-GNN, differing only in the local updates performed via hash functions.
@@
$\min~{}~{}\begin{bmatrix}~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1}}&~{}~{}~{}~{}~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{2}}&~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{3}}~{}~{}~{}\end{bmatrix}\begin{bmatrix}1\\ 2\\ 3\end{bmatrix}$s.t.
$2~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1}}$ $+$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{2}}$ $\leq 5$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{2}}$ $+$ $3~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1}}~{}~{}~{},~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{2}}~{}~{}~{}~{}~{},~{}~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{3}}~{}\leq 1$
$x_{1}\in\mathbb{Z}$
@@
With the concepts of color refinement and partition, we can introduce the core concept of this paper: message-passing-tractability.
@@
The proof of Theorem 4.4 is deferred to Appendix A. Let us compare our findings with [chen2022representing-milp] that establishes the existence of an MP-GNN capable of directly mapping $G$ to one of its optimal solutions, under the assumption that $G$ must be unfoldable. Unfoldability means that, after enough rounds of the WL test, each node receives a distinct color assignment. Essentially, it assumes that the WL test can differentiate between all nodes in $G$, and the elements within the corresponding partition $(\mathcal{I},\mathcal{J})$ have cardinality one: $|I_{p}|=1$ and $|J_{q}|=1$ for all $p\in\{1,2,\dots,s\}$ and $q\in\{1,2,\dots,t\}$. Consequently, any unfoldable MILP must be MP-tractable because the submatrices under the partition of an unfoldable MILP $(A_{ij})_{i\in I_{p},j\in J_{q}}$ must be $1\times 1$ and obviously satisfy the condition in Definition 4.2. However, the reverse assertion is not true: The example in Figure 2 serves as a case in point—it is MP-tractable but not unfoldable. Therefore, unfoldability is a stronger assumption than MP-tractability. Our Theorem 4.4 demonstrates that, to illustrate the expressive power of MP-GNNs in approximating SB, MP-tractability suffices; we do not need to make assumptions as strong as those required when considering MP-GNN for approximating the optimal solution.
Additionally, Theorem 4.4 offers practical insights. Before attempting to train a GNN to imitate SB, practitioners can initially verify if MILPs in their dataset satisfy MP-tractability. Notably, assessing MP-tractability relies solely on conducting the WL test (Algorithm 1). This algorithm is well-established in graph theory and benefits from abundant resources and repositories for implementation. Moreover, it operates with polynomial complexity, which is reasonable compared to solving MILPs.
@@ 4.2. MP-GNNs cannot universally represent SB beyond MP-tractability
The proof of Theorem 4.7 relies on the second-order folklore Weisfeiler-Lehman (2-FWL) test. Further details can be found in Appendix C.

### v3

@@ 1. Introduction
Branching is the process of dividing a linear relaxation into two subproblems. When branching, the solver selects a variable with a fractional value in the relaxation’s solution and create two new subproblems. In one subproblem, the variable is forced to be less than or equal to the nearest integer below the fractional value. In the other, it is bounded above the fractional value. The branching variable choice is critical because it can impact the solver’s efficiency by orders of magnitude.
@@
Learning to branch (L2B) introduces a new approach by incorporating machine learning (ML) to develop branching strategies, offering new solutions to address this trade-off. This line of research begins with imitation learning [khalil2016learning, alvarez2017machine, balcan2018learning, gasse2019exact, gupta2020hybrid, zarpellon2021parameterizing, gupta2022lookback, lin2022learning, yang2022learning], where models, including SVM, decision tree, and neural networks, are trained to mimic SB outcomes based on the features of the underlying MILP. They aim to create a computationally efficient strategy that achieves the effectiveness of SB on specific datasets. Furthermore, in recent reinforcement learning approaches, mimicking SB continues to take crucial roles in initialization or regularization [qu2022improved, zhang2022deep].
@@
  • In the context of L2B using GNNs, we first focus on the most widely used type: message-passing GNNs (MP-GNNs). Our study reveals that MP-GNNs can reliably predict SB results, but only for a specific class of MILPs that we introduce as message-passing-tractable (MP-tractable). We prove that for any distribution of MP-tractable MILPs, there exists an MP-GNN capable of accurately predicting their SB results. This finding establishes a theoretical basis for the widespread use of MP-GNNs to approximate SB results in current research.
@@ 2. Preliminaries and problem setup
Graph Representation of MILP
Here we present an approach to represent MILP as a bipartite graph, termed the MILP-graph. This conceptualization was initially proposed by [gasse2019exact] and has quickly become a prevalent model in ML for MILP-related tasks. The MILP-graph is defined as a tuple $G=(V,W,A,F_{V},F_{W})$, where the components are specified as follows: $V=\{1,2,\dots,m\}$ and $W=\{1,2,\dots,n\}$ are sets of nodes representing the constraints and variables, respectively. An edge $(i,j)$ connects node $i\in V$ to node $j\in W$ if the corresponding entry $A_{ij}$ in the coefficient matrix of (2.1) is non-zero, with $A_{ij}$ serving as the edge weight. $F_{V}$ are features/attributes of constraints, with features $v_{i}=(b_{i},\circ_{i})$ attached to node $i\in V$. $F_{W}$ are features/attributes of variables, with features $w_{j}=(c_{j},\ell_{j},u_{j},\delta_{I}(j))$ attached to node $j\in W$, where $\delta_{I}(j)\in\{0,1\}$ indicates whether the variable $x_{j}$ is integer-constrained.
[…]
@@
Message-passing graph neural networks (MP-GNNs) are a class of GNNs that operate on graph-structured data, by passing messages between nodes in a graph to aggregate information from their local neighborhoods. In our context, the input is an aforementioned MILP-graph $G=(V,W,A,F_{V},F_{W})$, and each node in $W$ is associated with a real-number output. We use the standard MP-GNNs for MILPs in the literature [gasse2019exact, chen2022representing-milp].
Specifically, the initial layer assigns features $s_{i}^{0},t_{j}^{0}$ for each node as
@@
Any $2$-FGNN, $F\in\mathcal{F}_{\textup{2-FGNN}}$, maps a MILP-graph $G$ to a $n$-dim vector: $y=F(G)$. While MP-GNNs and 2-FGNNs share the same input-output structure, their internal structures differ, leading to distinct expressive powers.
@@ 3. Imitating strong branching by GNNs
Observation I
SB is notably effective in reducing the size of the BnB search space. This size is measured by the size of the BnB tree. Here, a “tree” refers to a hierarchical structure of “nodes”, each representing a decision point or a subdivision of the problem. The tree’s size corresponds to the number of these nodes. For instance, consider the instance “neos-3761878-oglio” from MIPLIB [gleixner2021miplib]. When solved using SCIP [BolusaniEtal2024OO, BolusaniEtal2024ZR] under standard configurations, the BnB tree size is $851$, and it takes $61.04$ seconds to attain optimality. However, disabling SB, along with all branching rules dependent on SB, results in an increased BnB tree size to $35548$ and an increased runtime to $531.0$ seconds.
Observation II
SB itself is computationally expensive. In the above experiment under standard settings, SB consumes an average of $70.40\%$ of the total runtime, $42.97$ out of $61.04$ seconds in total.
@@
Remark: LP solution with the smallest $\ell_{2}$-norm
We only define the SB score for MILP problems with feasible and bounded LP relaxations; otherwise the optimal solution $x^{*}_{\textup{LP}}(G)$ does not exist. If the LP relaxation of $G$ admits multiple optimal solutions, then the strong branching score $\textup{SB}(G)$ depends on the choice of the particular optimal solution. To guarantee that the SB score is uniquely defined, in Definition 3.2, we use the optimal solution with the smallest $\ell_{2}$-norm, which is unique.
Remark: SB at leaf nodes
@@
Remark: Other types of SB scores
Although this paper primarily focuses on the product SB scores (where the SB score is defined as the product of objective value changes when branching up and down), our analysis can extend to other forms of SB scores in [dey2024theoretical]. (Refer to Appendix D.1)
@@ 4.1. MP-GNNs can represent SB for MP-tractable MILPs
In this subsection, we define a class of MILPs, named message-passing-tractable (MP-tractable) MILPs, and then show that MP-GNNs can represent SB within this class.
To define MP-tractability, we first present the Weisfeiler-Lehman (WL) test [weisfeiler1968reduction], a well-known criterion for assessing the expressive power of MP-GNNs [xu2019powerful]. The WL test in the context of MILP-graphs is stated in Algorithm 1. It follows exactly the same updating rule as the MP-GNN, differing only in the local updates performed via hash functions.
@@
With the concepts of color refinement and partition, we can introduce the core concept of this paper:
@@
The proof of Theorem 4.4 is deferred to Appendix A, with key ideas outlined here. First, we show that if Algorithm 1 produces identical results for two MP-tractable MILPs, they must share the same SB score. That is, if two MP-tractable MILPs have different SB scores, the WL test (or equivalently MP-GNNs) can capture this distinction. Building on this result, along with a generalized version of the Stone-Weierstrass theorem and Luzin’s theorem, we reach the final conclusion.
Let us compare our findings with [chen2022representing-milp] that establishes the existence of an MP-GNN capable of directly mapping $G$ to one of its optimal solutions, under the assumption that $G$ must be unfoldable. Unfoldability means that, after enough rounds of the WL test, each node receives a distinct color assignment. Essentially, it assumes that the WL test can differentiate between all nodes in $G$, and the elements within the corresponding partition $(\mathcal{I},\mathcal{J})$ have cardinality one: $|I_{p}|=1$ and $|J_{q}|=1$ for all $p\in\{1,2,\dots,s\}$ and $q\in\{1,2,\dots,t\}$. Consequently, any unfoldable MILP must be MP-tractable because the submatrices under the partition of an unfoldable MILP $(A_{ij})_{i\in I_{p},j\in J_{q}}$ must be $1\times 1$ and obviously satisfy the condition in Definition 4.2. However, the reverse assertion is not true: The example in Figure 2 serves as a case in point—it is MP-tractable but not unfoldable. Therefore, unfoldability is a stronger assumption than MP-tractability. Our Theorem 4.4 demonstrates that, to illustrate the expressive power of MP-GNNs in approximating SB, MP-tractability suffices; we do not need to make assumptions as strong as those required when considering MP-GNN for approximating the optimal solution.
@@ 4.2. MP-GNNs cannot universally represent SB beyond MP-tractability
The proof of Theorem 4.7 leverages the second-order folklore Weisfeiler-Lehman (2-FWL) test. We show that for any two MILPs, whether MP-tractable or not, identical 2-FWL results imply they share the same SB score, thus removing the need for MP-tractability. Details are provided in Appendix C.
@@
4.4. Practical insights of our theoretical results
Theorem 4.4 and Corollary 4.6 indicate the significance of MP-tractability in practice. Before attempting to train a MP-GNN to imitate SB, practitioners can first verify if the MILPs in their dataset satisfy MP-tractability. If the dataset contains a substantial number of MP-intractable instances, careful consideration of this approach is necessary, and 2-FGNNs may be more suitable according to Theorem 4.7. Notably, assessing MP-tractability relies solely on conducting the WL test (Algorithm 1). This algorithm is well-established in graph theory and benefits from abundant resources and repositories for implementation. Moreover, it operates with polynomial complexity (detailed below), which is reasonable compared to solving MILPs.
Complexity of verifying MP-tractability
To verify MP-tractability of a MILP, one requires at most $\mathcal{O}(m+n)$ color refinement iterations according to Theorem 4.1. The complexity of each iteration is bounded by the number of edges in the graph [shervashidze2011weisfeiler]. In our context, it is bounded by the number of nonzeros in matrix $A$: $\text{nnz}(A)$. Therefore, the overall complexity is $\mathcal{O}((m+n)\cdot\text{nnz}(A))$, which is linear in terms of $(m+n)$ and $\text{nnz}(A)$. In contrast, solving an MILP or even calculating its all the SB scores requires significantly higher complexity. To calculate the SB score of each MILP, one needs to solve at most $n$ LPs. We denote the complexity of solving each LP as $\text{CompLP}(m,n)$. Therefore, the overall complexity of calculating SB scores is $\mathcal{O}(n\cdot\text{CompLP}(m,n))$. Note that, currently, there is still no strongly polynomial-time algorithm for LP, thus this complexity is significantly higher than that of verifying MP-tractability.
While verifying MP-tractability is polynomial in complexity, the complexity of GNNs is still not guaranteed. Theorems 4.4 and 4.7 address existence, not complexity. In other words, this paper answers the question of whether GNNs can represent the SB score. To explore how well GNNs can represent SB, further investigation is needed.
Frequency of MP-tractability
In practice, the occurrence of MP-tractable instances is highly dependent on the dataset. In both Examples 4.1 and 4.2 (both MP-intractable), all variables exhibit symmetry, as they are assigned the same color by the WL test, which fails to distinguish them. Conversely, in the 3-variable example in Figure 2 (MP-tractable), only two of the three variables, $x_{1}$ and $x_{2}$, are symmetric. Generally, the frequency of MP-tractability depends on the level of symmetry in the data — higher levels of symmetry increase the risk of MP-intractability. This phenomenon is commonly seen in practical MILP datasets, such as MIPLIB 2017 [gleixner2021miplib]. According to [chen2022representing-milp], approximately one-quarter of examples show significant symmetry in over half of the variables.